His 13- foot ladder leans against the building at a point that is 11.3 feet above the ground. 

At what height would a 15-foot ladder touch the building if both ladders form the same angle with the ground? Round your answer to the nearest tenth.

 ____________ feet 



1
Expert's answer
2021-12-16T17:10:51-0500



Angle formed by the 13- foot ladder and the ground is "sin^{-1}(11.3\/13)=60.4^0"

The height a 15-foot ladder would touch the building, "h\/15=sin60.4^0"

"h=15\times sin60.4^0=13.0feet"